# Nearly-zero large-angle anisotropy of the cosmic microwave background

### v5

@@ 4.2.4 Rank comparison with standard realizations
The middle panel shows the same quantities evaluated for $\Delta_{\text{minimal}}$, with the dipole-term adjustment described above. We again find that a small fraction of standard model realizations, ranging from about $10^{-1.8}$ (for WMAP) to $10^{-2.8}$ (for Commander), come as close to zero $\Delta_{\text{minimal}}$ as the measured CMB temperature maps. The values of $C_{1}$ for the best-fit dipole for the maps NILC, SMICA, COMMANDER, WMAP without its monopole, and WMAP with its monopole are approximately $365$, $341$, $322$, $392$, and $426$ $\mu$K2, respectively.
The lower panel shows the same quantities evaluated for $\Delta_{\text{maximal}}$, with the dipole-term adjustment described above. We again find that a small fraction of standard model realizations, ranging from about $10^{-2.3}$ (for WMAP) to $10^{-3.2}$ (for Commander), come as close to zero $\Delta_{\text{maximal}}$ as the measured CMB temperature maps. The values of $C_{1}$ for the best-fit dipole for the maps NILC, SMICA, COMMANDER, WMAP without its monopole, and WMAP with its monopole are approximately $384$, $390$, $389$, $440$, and $471$ $\mu$K2, respectively.
@@
5 Interpretation
In standard inflation, correlation varies randomly among different realizations, and the tiny measured correlation must be interpreted as a statistical anomaly. Our rank comparison (Fig. 4) shows that the sky agrees with zero better than almost all standard realizations. The most direct comparison, as well as the smallest measured values of correlation, appear in even-parity correlation, where Planck maps show deviations from zero that are three to four orders of magnitude smaller than typical standard realizations. The standard interpretation is that this small correlation happens by chance, which is estimated to occur in a fraction $10^{-2.8}$ to $10^{-4.3}$ of realizations, depending on the map.
Another interpretation is that the nearly-zero correlation is due to initial conditions more symmetric than generally assumed. One possibility is an exact fundamental causal symmetry of quantum gravity that is not included in the standard model. The measured angular range of minimal correlation agrees with that derived here from a basic physical principle of causal coherence— that quantum fluctuations generate physical correlations only within causal diamonds bounded by inflationary horizons. In this interpretation, the small measured departure from zero correlation is attributed to measurement error, dominated by contamination by the Galaxy. This view is consistent with measured variation among the different maps.
The measurements appear to be consistent with the interpretation that CMB correlation on large angular scales preserves a signature of a new causal symmetry of initial conditions. The symmetry is not consistent with the standard effective quantum field theory approximation for formation of inflationary perturbations, but could arise from a deeper structure of quantum geometry that becomes significant on the scale of causal horizons.
@@ 6 Conclusion
The surprisingly small absolute value of the large-angle CMB correlation function has been known since the first measurements with COBE (Bennett et al., 1994; Hinshaw et al., 1996). Subsequent measurements from WMAP (Bennett et al., 2003, 2011) and then Planck (Planck Collaboration, 2020c) show values successively closer to zero. This anomaly and others have been thoroughly analyzed (e.g., (Copi et al., 2009; Muir et al., 2018)), but are not generally thought to present a compelling challenge to standard cosmological theory, both because the standard theory occasionally produces such small correlations by chance, and because there has not been a precisely formulated and physically compelling alternative expectation.
@@
The significant new fact reported in this paper is that when the removed dipole component is accounted for, the magnitude of correlations is much smaller than previously documented. In particular, direct measurements of even-parity correlation in the Planck data at angular separations in the geometrically-calculated causal shadow are consistent with zero, with measured dimensionless fractional variations smaller than $10^{-25}$. Correlations are several orders of magnitude smaller than expected in standard cosmology, even though it starts with a perfectly isotropic unperturbed classical background; depending on the map, the probability of such small correlations occurring by chance in the Planck maps is $10^{-2.8}$ to $10^{-4.3}$.
@@
We are thus led to suspect that nearly-zero large angle correlation may not be an extraordinary anomaly of our particular sky, but a signature of new physical symmetry in cosmic initial conditions, which are shaped by quantum gravity. As one example, we proposed a geometrically-defined symmetry that invokes causal, nonlocal “spooky” quantum entanglement of gravitational vacuum fluctuation states, which is not included in the standard theory of inflationary perturbations based on quantum field theory, but is a plausible feature for a deeper theory of quantum gravity that addresses entanglement with causal horizons. As discussed in the Appendix, a symmetry of this kind is broadly consistent with tests of classical concordance cosmology, which mainly (if not entirely) depend only on a nearly scale-invariant initial 3D power spectrum of perturbations averaged over large volumes. It is also consistent with all other experimental tests of QFT, none of which depend on quantized gravity.
@@
If the causal-symmetry hypothesis is not true, it can be falsified by more precise measurement of nonzero correlations on the cosmic horizon within predicted causal shadows. The precision of the results reported here, and the significance of our null tests, are not limited by any fundamental source of noise, but by the accuracy of models of Galactic foreground emission. Tests of the symmetry can be improved with all-sky models of emission from the Galaxy that allow better measurements of the true CMB pattern on the largest scales. In this analysis, we have used the published all-sky, unmasked, Galaxy-subtracted CMB maps prepared by the satellite teams. We have not undertaken a close comparative study of the Galaxy models used to create those maps, but it is plausible that existing data from WMAP and Planck could be combined to create maps better optimized for studies of large-angle correlation symmetry. Galaxy models would improve with better spectral measurements, as proposed in new satellite concepts (Chluba et al., 2021). Other unique signatures of causal coherence are addressed briefly in the Appendix.

### v7

@@ 5.2.4 Rank comparison with standard realizations
[…]
@@
5.3 Interpretation
In standard inflation, correlation varies widely among different realizations, and the tiny measured correlation must be interpreted as a statistical anomaly. Our rank comparison (Fig. 4) shows that the sky agrees with zero better than almost all standard realizations. The most direct comparison, as well as the smallest measured values of correlation, appear in even-parity correlation.
Another interpretation is that the nearly-zero correlation is due to initial conditions more symmetric than generally assumed. One possibility is an exact fundamental causal symmetry that is not included in the standard model. The measured angular range of minimal correlation agrees with a range of zero correlation derived here from 4D causal coherence. In this interpretation, the measured departures from zero correlation are attributed to measurement error, dominated by contamination by the Galaxy. This view is consistent with measured variation among the different maps.
@@ 6 Conclusion
A surprisingly small absolute value of the large-angle CMB correlation function has been known since the first measurements with COBE. Subsequent measurements from WMAP and then Planck showed values successively closer to zero. They are not generally thought to present a compelling challenge to standard cosmological theory, both because the standard theory occasionally produces such small correlations by chance, and because there has not been a precisely formulated and physically compelling alternative expectation.
@@
The significant new fact reported in this paper is that when the removed dipole component is accounted for, the magnitude of directly measured correlation over a range of angles around $\Theta=90^{\circ}$ is much smaller than previously documented. Allowing for measurement uncertainty, even-parity correlation in a geometrically-calculated range of angles is consistent with zero, and at least several orders of magnitude smaller than expected from standard initial conditions. The probability of correlations as small as those measured in Planck maps is $10^{-2.8}$ to $10^{-4.3}$.
We are thus led to suspect that nearly-zero large angle correlation may not be an accident of our particular sky, but a signature of physical symmetry in cosmic initial conditions. A range of zero correlation that matches the data can be calculated geometrically from a causal bound on the coherence of gravitational vacuum fluctuation states, which is not included in the standard theory of inflationary perturbations based on QFT.
@@
As discussed in the Appendix, a symmetry of this kind is broadly consistent with tests of classical concordance cosmology, which mainly (if not entirely) depend only on a nearly scale-invariant initial 3D power spectrum of perturbations averaged over large volumes. It is also consistent with all other experimental tests of QFT, none of which depend on quantized gravity.
@@
If the causal-symmetry hypothesis is not true, it can be falsified by more precise measurement of nonzero correlations on the cosmic horizon within predicted causal shadows. The precision of the results reported here, and the significance of our null tests, are not limited by any fundamental source of noise, but by the accuracy of models of Galactic foreground emission. Tests of the symmetry could be improved with all-sky models of emission from the Galaxy that allow more accurate measurement of the true CMB pattern on the largest scales. Other unique cosmological signatures of causal coherence are addressed briefly in the Appendix.